HIGH RESOLUTION VELOCITY-MAP IMAGED PHOTODETACHMENT SPECTRA OF O^{−} AND OH^{−}

\underline{S. T. GIBSON}, S. J. CAVANAGH, M. N. GALE, E. H. ROBERTS, and B. R. LEWIS, Research School of Physical Sciences and Engineering, The Australian National University, Canberra ACT 0200, Australia (email: Stephen.Gibson@anu.edu.au); J. R. GASCOOKE, School of Chemistry and Physics, The University of Adelaide, Adelaide SA 5005, Australia.

The 532 nm photodetachment spectra of O^{−} and OH^{−} have been measured using a negative-ion beam spectrometer which incorporates a velocity-map imaging (VMI) lens,^{a} located co-axially within the ion beam. The VMI technique offers a number of advantages for the recording of photoelectron spectra, including the simultaneous detection of all kinetic energies and the complete angular distribution of photoelectrons. To date, the energy-resolution reported for photoelectron spectra using imaging techniques have been limited to ΔE/E ≥ 2%^{b} with improvement in energy-resolution achieved only through the use of slow electrons. Our spectrum for O^{−} was recorded with electron energies near 0.87 eV with ΔE/E ≤ 0.5%. This is a significant achievement for this technique, providing spectra with considerable detail, where individual fine-structure and some rotational transitions are resolved. Measurements above threshold provide more stringent tests on the usefulness of near threshold theories of photodetachment. The VMI image of O^{−} is visually similar to OH^{−}, with a propensity for the electron to be ejected at 90^{∘} to the laser polarization (asymmetry parameter β ∼ −1), reflecting the detachment of an electron from the p−orbital of the oxygen atom. Detail in the spectra reveal different angular distributions for individual transitions, reflecting the nature of the fine-structure transition and the interference between partial waves.

aA. T. J. B. Eppink and D. H. Parker, Rev. Sci. Instrum. 68, 3477 (1997).

bS. M. Sheehan, G. Meloni, B. F. Parsons, N. Wehres, and D. M. Neumark, J. Chem. Phys. 124, 064303 (2006).